7. A company had net sales of P 600,00 during the past year. it's inventory of goods was the same quantity at the beginning and at the end of year. It's cost of goods sold consisted of $ 120,000 of variable costs and $ 200,000 of fixed costs. Its selling and administrative expenses were $ 40,000 of variable and $ 150,000 of fixed expenses. Find the company's gross margin?









8.Company ABC makes shoes. If ABC reported $ 5,000.000 in total revenue for the year and cost of goods sold ( cost of materials and direct labor ) of $ 2.5 million, then we can use the formula above to find ABC's gross profit margin?







9. Your lemonade stand earned P 5,000 but it cost you P 4,000 to buy lemons. What is your groos profit margin?









10. A service firm that charges $ 1,000 for services and pays its employee $ 600 to deliver those services earns a gross margin of _______sales.
1
Expert's answer
2021-09-13T17:10:30-0400

7)

"Gross\space margin=\frac{net\space sales-cost\space of \space goods}{net \space sales}\\=\frac{6000-(120000+200000+40000+150000)}{600000}\\=15%"

8)

"Gross\space profit\space margin=\frac{total\space revenue-cost\space of \space goods}{total\space revenue}\\=\frac{5000000-2500000}{5000000}\\=50%"

9)

"Gross\space profit\space margin=\frac{total\space revenue-cost\space of \space goods}{total\space revenue}\\=\frac{5000-4000}{5000}\\=20%"

10)

"Gross\space margin=\frac{total\space revenue-cost\space of \space goods}{total\space revenue}\\=\frac{1000-600}{1000}\\=40%"